OpenStudy (anonymous):

Question 3: Ellen works for a high-speed rail company that wants to develop a new rail line. Ellen’s project is to find a train that is the second fastest in the world. • The Shinkansen Bullet Train in Japan is reported to go as fast as 320 kilometers per hour. • The TGV train in France can reach speeds of 89.44 meters per second. Explain to Ellen how to find what speed her new train must go to be the second fastest in this group. Then, find an appropriate speed for Ellen’s train in miles per hour and use the formula distance = speed • time to find how far the train can go in 2.5 hours

OpenStudy (anonymous):

Well to find the speed of the TGV train, you know that 1 km = 1000 meters; 1 hour = 3600 seconds. So with 89.44 m/s you can do,\[\frac{ 89.44 meters }{ second }*\frac{ 1 km }{ 1000meters }*\frac{ 3600seconds }{ 1hour }=\frac{ ?km }{ ?hr }\]The way how I setup the proportions is I made sure that I cancelled the older units (m/s) to get to the newer units (km/hr).

OpenStudy (anonymous):

thankss @hwtrain

OpenStudy (anonymous):

For the final part, you know that 1mile = roughly 1.609 km. So, once you get the answer from the other part (in km/hr) make sure to times by (1 mile/1.609km) to get from km to miles. Then you would multiply by 2.5.

OpenStudy (anonymous):

ok :)